Solve this:
P²(x²-a²)-2pxy+y²-b²=0
1
Expert's answer
2020-12-20T18:41:02-0500
"p^2(x^2-a^2)-2pxy+y^2-b^2=0"

"p^2x^2-p^2a^2-2pxy+y^2-b^2=0"

"(y^2-2pxy+p^2x^2)-(p^2a^2+b^2)=0"

"(y-px)^2=p^2a^2+b^2"

"y=px\pm\sqrt{p^2a^2+b^2}"

Each of the component equations is in Clairaut's form.

Hence changing "p" to the arbitrary constant "c" in


"p^2(x^2-a^2)-2pxy+y^2-b^2=0"


the required solution is


"c^2(x^2-a^2)-2cxy+y^2-b^2=0"

Or


"y=cx\pm\sqrt{c^2a^2+b^2}, c=const"